# Residual permutation test for regression coefficient testing

### v3

@@ 4 Residual permutation test: methodology and size validity
In Section 3, we have shown from simulation experiments that a naive permutation test on the residuals, although more robust than ANOVA, is still not guaranteed to have finite-population size validity with just exchangeable noise. In this section we describe a more refined test using the projected residuals $\hat{\boldsymbol{\varepsilon}}$ and $\hat{\boldsymbol{e}}$, which we call the residual permutation test (RPT), and present its finite-population size validity guarantee in Theorem 2. For intuition behind such construction, we refer the readers to Section 4.1. We will assume throughout this section that the design matrix $(\boldsymbol{X},\boldsymbol{Z})$ is deterministic.
@@
We remark that as shown in Theorem 2, an important advantage of RPT is that the result is finite-population such that it holds for arbitrary size of $n$. Moreover, our result assumes a fixed-design matrix and does not require any assumption on $\boldsymbol{X}$ for finite-population size validity. For example, the rank of $\boldsymbol{X}$ even does not necessarily need to be $p$. Also, Theorem 2 shows that RPT has valid size for any choice of function $T(\cdot,\cdot)$ and number of permutations $K$. However, in practice, to have good power under the alternative, we typically set $T(x,y)=|\langle x,y\rangle|$ and choose a moderate size of $K=O(1/\alpha)$.
@@ 4.1 Some intuition of RPT
In this section, we discuss the intuition behind (5). As demonstrated in Section 3, a naive permutation test on the residuals is in general not valid in the finite population setting with just exchangeable noises. This is because under the null, $\phi_{\mathrm{naive}}$ performs permutations on the vector $\hat{\boldsymbol{\varepsilon}}=\boldsymbol{V}_{0}^{\top}\boldsymbol{\varepsilon}$ instead of $\boldsymbol{\varepsilon}$ itself. Even if $\boldsymbol{\varepsilon}$ is an exchangeable random vector, $\boldsymbol{V}_{0}^{\top}\boldsymbol{\varepsilon}$ may no longer be so, which renders the naive test invalid.
@@ 5 Analysis of statistical power
This section provides power analysis of RPT under mild moment assumptions of noises $\varepsilon_{i}$ and $e_{i}$’s where, e.g., the second order moments are not necessarily finite. For simplicity of exposition, throughout this section we assume without loss of generality that $n$ is a multiple of $|\mathcal{P}_{K}|=K+1$, where $K$ is a fixed constant that is chosen such that $K\geq 1/\alpha$ for the prespecified Type-I error $\alpha$. The scenario where $n$ is not divisible by $K+1$ can be handled by randomly discarding a subset of data of size at most $K$ to make $n$ divisible. We will focus on the version of RPT defined in (6) with $T(x,y)=|\langle x,y\rangle|$. While we continue to assume that the design matrix $\boldsymbol{X}$ is deterministic, $\boldsymbol{Z}$ is assumed to have a random design following specific models in this and the next two sections. Moreover, we are primarily interested in the dependence of the power of RPT on the tail heaviness of the noise distributions. To this end, we make the following assumption on the model:
@@
The statistical power guarantee in Theorem 3 requires the set of permutations to follow Assumption 4, whilst the finite-population size validity requires instead Assumption 2. Then an important question is, how to effectively construct a $\mathcal{P}_{K}$ that satisfies both assumptions. In Section 5.1, we provide an algorithm to answer this question. In order to prove Theorems 3 and 4, we are faced with two questions, the first is that we do not have any assumption on $\boldsymbol{X}$, so that $\tilde{\boldsymbol{V}}_{j}$ can follow arbitrary pattern; the second is the heavy tails of $e_{i}$’s and $\varepsilon_{i}$’s. We defer the proof of the two theorems to the Supplementary Material. To help the readers understand the intuitions of the proof, we provide a proof sketch of the main Theorem 3 in Section 5.2.
@@ 3       Generate an independent random permutation $\pi$ of indices $\{1,\ldots,n\}$
Notice that throughout this article, we assume that the alternative class is in the form $\boldsymbol{Y}=\boldsymbol{X}\beta+b\boldsymbol{Z}+\boldsymbol{\varepsilon}$ for some $b\neq 0$, whence we invoke Assumption 4 to increase its statistical power. When the alternative class follows other forms, such as $\boldsymbol{Y}=\boldsymbol{X}\beta+f(\boldsymbol{Z})+\boldsymbol{\varepsilon}$ with some nonlinear function $f:\mathbb{R}^{n}\mapsto\mathbb{R}^{n}$, one may not necessarily need Assumption 4 anymore. Instead, one may need other assumptions on $\mathcal{P}_{K}$ to adapt to the nonlinear function $f(\cdot)$. In light of Algorithm 1 and our theoretical statements, we summarize an implementation of RPT in Algorithm 2. The maximum time complexities of Algorithms 1 and 2 are $O(TKnp)$ and $O(TKnp+Knp^{2})$ respectively, where $T$ is the maximum number of iterations. The expected time complexities of the two algorithms are instead $O(Knp)$ and $O(Knp^{2})$, respectively.
@@ 7 Minimax rate optimality of coefficient tests
$\mathcal{R}_{t,\boldsymbol{X}}(\tau):=\inf_{\varphi\in\Phi}\left\{\sup_{\mathbb{P}_{\varepsilon}\in\mathcal{D}_{t}}\sup_{\mathbb{P}_{e}\in\mathcal{D}_{1}\cap\tilde{\mathcal{D}}}\sup_{\beta,\beta^{Z}\in\mathbb{R}^{p}}\mathbb{P}_{0}(\varphi=1)+\sup_{|b|\geq\tau}\sup_{\mathbb{P}_{\varepsilon}\in\mathcal{D}_{t}}\sup_{\mathbb{P}_{e}\in\mathcal{D}_{1}\cap\tilde{\mathcal{D}}}\sup_{\beta,\beta^{Z}\in\mathbb{R}^{p}}\mathbb{P}_{b}(\varphi=0)\right\}.$
@@ 8.3 Numeric analysis of alternative power
From Figures 2(a)-(c), (d) and (f), we can conclude that in most of the settings, the power of RPT is slightly worse than ANOVA, the naive RPT and FL. The difference is more pronounced when both the design and the noise follow a heavy-tailed distribution (Figure 2(e)). However, bearing in mind the lack of valid size control of ANOVA, naive RPT, DR, FL and CRT, especially when design and noise are heavy-tailed, we would argue that the gap in power between RPT and these competitors is the price to pay for finite-population size validity with only exchangeable noise in the proportional regime. Moreover, RPT is nevertheless still guaranteed to reject the alternative with high probability given a signal size $b$ not too much larger than the competitors. In addition, we observe that DR does not seem to have power converging to 1 as $b$ increases for heavy-tailed noise, while the power of CRT is substantially reduced for heavy-tailed design distributions.
@@ 9 Discussion
In this paper, we propose a new method for fixed design regression coefficient test when the number of covariates $p$ can be as large as a fraction of the sample size $n$. RPT is a permutation-based approach that exploits the exchangeability of the noise terms to achieve finite-population size control. Our approach uses the fact that the empirical residuals of the classical OLS fit is equivalent to the projection of the noise vector onto an subspace orthogonal to the design to construct a test with valid size for $p<n/2$ based on multiple subspace projection. At the same time, we provide power analysis of RPT, and derived the signal detection rate of the coefficient $b$ in the presence of heavy-tailed noise vector $\boldsymbol{\varepsilon}$. As a by product, we propose $\textrm{RPT}_{\textrm{EM}}$ and demonstrate its size validity and power via numerical experiments. It would be of interest to understand the theoretical properties of $\textrm{RPT}_{\textrm{EM}}$ in future study.
In the regime where $n/2\leq p<n$, we propose the naive RPT, and prove its finite-population size validity under spherically invariant distributions, and compare it with ANOVA as well as other competing approaches via numerical experiments. In the meanwhile, we provide a more profound analysis of ANOVA test, which is of independent interest for practitioners interested in ANOVA.
@@
A2 Proof of finite-population size validity statements

### v4

@@ 4 Residual permutation test: methodology and size validity
In Section 3, we have shown from simulation experiments that a naive permutation test on the residuals, although more robust than ANOVA, is still not guaranteed to have finite-sample size validity with just exchangeable noise. In this section we describe a more refined test using the projected residuals $\hat{\boldsymbol{\varepsilon}}$ and $\hat{\boldsymbol{e}}$, which we call the residual permutation test (RPT), and present its finite-sample size validity guarantee in Theorem 2. For intuition behind such construction, we refer the readers to Section 4.1. We will assume throughout this section that the design matrix $(\boldsymbol{X},\boldsymbol{Z})$ is deterministic.
@@
We remark that as shown in Theorem 2, an important advantage of RPT is that the result is finite sample in the sense that it holds for arbitrary size of $n$. Moreover, our result assumes a fixed-design matrix and does not require any assumption on $\boldsymbol{X}$ for finite-sample size validity. For example, the rank of $\boldsymbol{X}$ even does not necessarily need to be $p$. Also, Theorem 2 shows that RPT has valid size for any choice of function $T(\cdot,\cdot)$ and number of permutations $K$. However, in practice, to have good power under the alternative, we typically set $T(x,y)=|\langle x,y\rangle|$ and choose a moderate size of $K=O(1/\alpha)$.
@@ 4.1 Some intuition of RPT
In this section, we discuss the intuition behind (5). As demonstrated in Section 3, a naive permutation test on the residuals is in general not valid in the finite-sample setting with just exchangeable noises. This is because under the null, $\phi_{\mathrm{naive}}$ performs permutations on the vector $\hat{\boldsymbol{\varepsilon}}=\boldsymbol{V}_{0}^{\top}\boldsymbol{\varepsilon}$ instead of $\boldsymbol{\varepsilon}$ itself. Even if $\boldsymbol{\varepsilon}$ is an exchangeable random vector, $\boldsymbol{V}_{0}^{\top}\boldsymbol{\varepsilon}$ may no longer be so, which renders the naive test invalid.
@@ 5 Analysis of statistical power
This section provides power analysis of RPT under mild moment assumptions of noises $\varepsilon_{i}$ and $e_{i}$’s when the second order moments are not necessarily finite. For simplicity of exposition, throughout this section we assume without loss of generality that $n$ is a multiple of $|\mathcal{P}_{K}|=K+1$, where $K$ is a fixed constant that is chosen such that $K\geq 1/\alpha$ for the prespecified Type-I error $\alpha$. The scenario where $n$ is not divisible by $K+1$ can be handled by randomly discarding a subset of data of size at most $K$ to make $n$ divisible. We will focus on the version of RPT defined in (6) with $T(x,y)=|\langle x,y\rangle|$. While we continue to assume that the design matrix $\boldsymbol{X}$ is deterministic, $\boldsymbol{Z}$ is assumed to have a random design following specific models in this and the next two sections. Moreover, we are primarily interested in the dependence of the power of RPT on the tail heaviness of the noise distributions. To this end, we make the following assumption on the model:
@@
The statistical power guarantee in Theorem 3 requires the set of permutations to follow Assumption 4, whilst the finite-sample size validity requires instead Assumption 2. Then an important question is, how to effectively construct a $\mathcal{P}_{K}$ that satisfies both assumptions. In Section 5.1, we provide an algorithm to answer this question. In order to prove Theorems 3 and 4, we are faced with two questions, the first is that we do not have any assumption on $\boldsymbol{X}$, so that $\tilde{\boldsymbol{V}}_{j}$ can follow arbitrary pattern; the second is the heavy tails of $e_{i}$’s and $\varepsilon_{i}$’s. We defer the proof of the two theorems to the Supplementary Material. To help the readers understand the intuitions of the proof, we provide a proof sketch of the main Theorem 3 in Section 5.2.
@@ 3       Generate an independent random permutation $\pi$ of indices $\{1,\ldots,n\}$
Notice that throughout this article, we assume that the alternative class is in the form $\boldsymbol{Y}=\boldsymbol{X}\beta+b\boldsymbol{Z}+\boldsymbol{\varepsilon}$ for some $b\neq 0$, whence we invoke Assumption 4 to increase its statistical power. When the alternative class follows other forms, such as $\boldsymbol{Y}=\boldsymbol{X}\beta+f(\boldsymbol{Z})+\boldsymbol{\varepsilon}$ with some nonlinear function $f:\mathbb{R}^{n}\mapsto\mathbb{R}^{n}$, one may not necessarily need Assumption 4 anymore. Instead, one may need other assumptions on $\mathcal{P}_{K}$ to adapt to the nonlinear function $f(\cdot)$. In light of Algorithm 1 and our theoretical statements, we summarize an implementation of RPT in Algorithm 2. The maximum time complexities of Algorithms 1 and 2 are $O(TKnp)$ and $O(TKnp+Knp^{2})$ respectively, where $T$ is the maximum number of iterations. The expected time complexities of the two algorithms are instead $O(Knp)$ and $O(Knp^{2})$, respectively. We also remark that due to the construction of the permutation set in Algorithm 1, RPT is inherently a randomized procedure, and unlike permutation tests or bootstrapping, this variability due to randomness cannot be reduced by increasing computational time. It would be of interest to propose a new test that eradicates such reproducibility issue while maintaining all the beneficial features of Algorithm 1, which we leave for future work.
@@ 6 Statistical power under broader classes of alternatives
$\sum_{i=1}^{n}\mathbb{P}\left(|\varepsilon_{i}|^{1+t}\geq Bi\right)$ $\leq\sum_{i=1}^{n}\mathbb{P}\left(|\varepsilon_{\infty}|^{1+t}\geq Bi\right)$
@@ 7 Minimax rate optimality of coefficient tests
$\mathcal{R}_{t,\boldsymbol{X}}(\tau):=$
$\quad\inf_{\varphi\in\Phi}\left\{\sup_{\mathbb{P}_{\varepsilon}\in\mathcal{D}_{t}}\sup_{\mathbb{P}_{e}\in\mathcal{D}_{1}\cap\tilde{\mathcal{D}}}\sup_{\beta,\beta^{Z}\in\mathbb{R}^{p}}\mathbb{P}_{0}(\varphi=1)+\sup_{|b|\geq\tau}\sup_{\mathbb{P}_{\varepsilon}\in\mathcal{D}_{t}}\sup_{\mathbb{P}_{e}\in\mathcal{D}_{1}\cap\tilde{\mathcal{D}}}\sup_{\beta,\beta^{Z}\in\mathbb{R}^{p}}\mathbb{P}_{b}(\varphi=0)\right\}.$
@@ 8.3 Numeric analysis of alternative power
From Figures 2(a)-(c), (d) and (f), we can conclude that in most of the settings, the power of RPT is slightly worse than ANOVA, the naive RPT and FL. The difference is more pronounced when both the design and the noise follow a heavy-tailed distribution (Figure 2(e)). However, bearing in mind the lack of valid size control of ANOVA, naive RPT, DR, FL and CRT, especially when design and noise are heavy-tailed, we would argue that the gap in power between RPT and these competitors is the price to pay for finite-sample size validity with only exchangeable noise in the proportional regime. Moreover, RPT is nevertheless still guaranteed to reject the alternative with high probability given a signal size $b$ not too much larger than the competitors. In addition, we observe that DR does not seem to have power converging to 1 as $b$ increases for heavy-tailed noise, while the power of CRT is substantially reduced for heavy-tailed design distributions.
@@ 9 Discussion
In this paper, we propose a new method for fixed design regression coefficient test when the number of covariates $p$ can be as large as a fraction of the sample size $n$. RPT is a permutation-based approach that exploits the exchangeability of the noise terms to achieve finite-sample size control. Our approach uses the fact that the empirical residuals of the classical OLS fit is equivalent to the projection of the noise vector onto an subspace orthogonal to the design to construct a test with valid size for $p<n/2$ based on multiple subspace projection. At the same time, we provide power analysis of RPT, and derived the signal detection rate of the coefficient $b$ in the presence of heavy-tailed noise vector $\boldsymbol{\varepsilon}$. As a by product, we propose $\textrm{RPT}_{\textrm{EM}}$ and demonstrate its size validity and power via numerical experiments. It would be of interest to understand the theoretical properties of $\textrm{RPT}_{\textrm{EM}}$ in future study.
In the regime where $n/2\leq p<n$, we propose the naive RPT, and prove its finite-sample size validity under spherically invariant distributions, and compare it with ANOVA as well as other competing approaches via numerical experiments. In the meanwhile, we provide a more profound analysis of ANOVA test, which is of independent interest for practitioners interested in ANOVA.
@@
{acks}
[Acknowledgments] The authors would like to thank the Editor, the Associate Editor and two anonymous referees for helpful comments that improved the paper. YW is also affiliated with Shanghai Artificial Intelligence Laboratory and Shanghai Qi Zhi Institute. {funding} The research of TW was supported by EPSRC grant EP/T02772X/2. The research of YW was supported by National Key R & D Program (2022YFA1008100), and the grant of National Natural Science Foundation of China (NSFC) 12201341.
[…]
@@
A2 Proof of finite-sample size validity statements